Kim Distribution Services (KDS) distributes food purchased in bulk to small retailers. The firm is divided into two divisions: Purchasing and Distribution. Purchasing is responsible for ordering goods from the manufacturer, receiving them, and then moving them to the appropriate location in the warehouse. Distribution is responsible for taking orders from retailers, picking the products from the warehouse for the orders, and packaging the orders for shipment. KDS has a policy of filling every order on the day of the order. If the firm is out of a particular item, it will ship a partial order and complete the order when the item is back in stock. Occasionally, an order will not have to be packaged if the retailer chooses to take delivery at the KDS loading dock.

Distribution has 15 employees who are responsible for the activities, and all 15 are trained to handle any of the three tasks. Each of these employees works 40 hours per week for 50 weeks. There is an allowance of 15 percent of the employees’ time for training and other administrative tasks. The total costs of distribution for the coming year are estimated to be $948,600. When asked, the manager of Distribution estimated the following times for each of the three activities:

  • Taking orders: 10 minutes.

  • Picking orders: 14 minutes.

  • Packaging orders: 20 minutes.

During the year, the Distribution received 25,800 orders. Because of out-of-stock events, pickers had to pick 30,800 orders. 28,500 orders were packaged.

Required:
a. What is the cost per minute for activities in Distribution? (Round your answer to 2 decimal places.)

b. What is the cost of an order that requires all three activities? (Round your intermediate calculations and final answer to 2 decimal places.)

c. How many minutes of unused capacity did Distribution have for the year?

d. What was the cost of the unused capacity in Distribution? (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole dollar.)
Answer #1

SOLUTION

Total working minutes available = 15 * 40 * 50 * 60 minutes * (1-15%) = 1,530,000

Total actual working minutes used during the year = Taking orders (10m*25,800) + Picking orders (14m*30,800) + Packaging order (20m*28,500)

= 258,000 + 431,200 + 570,000 = 1,259,200

a. The cost per minute for activities in Distribution= Total distribution costs / Total working minutes available

= $948,600 / 1,530,000 = 0.62 per available minute

b. The cost of an order that requires all three activities = (10 + 14 + 20) * 0.62

= $27.28

c. Total minutes of unused capacity the Distribution have for the year = Available minutes - Used minutes

= 1,530,000 - 1,259,200 = 270,800 minutes

d. The cost of the unused capacity in Distribution = 270,800 * $0.62 = 167,896